### v1

# High-dimensional quantum XYZ product codes for biased noise

Abstract
Quantum XYZ product can construct a class of non-CSS codes by using three classical codes. However, before this work, their error-correcting performance is not studied in depth and whether this code construction method can be generalized to higher dimension is an open question. In this paper, we first study the error-correcting performance of the 3D Chamon code, which can be seen as a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes. Second, we show that XYZ product can be generalized to four dimension and propose four-dimensional (4D) XYZ product code construction, which can be seen as a variant of 4D homological product and constructs a class of non-CSS codes by using 4 classical codes or 2 CSS codes. Compared with 4D homological product, we show that 4D XYZ product can construct non-CSS codes with higher dimension or code distance. Third, we consider two special instances of 4D XYZ product, which we name 4D Chamon code and 4D XYZ concatenated code. Exploiting fully decoupled binary belief propagation combined with ordered statistics decoding, our simulation results show that, using the same two CSS codes, 4D XYZ product can construct non-CSS codes with better error-correcting performance for $Z$-biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
I Introduction
At present, even for the state-of-the-art quantum computer, qubits are still highly susceptible to noise in storage and computation, making it difficult to fulfil the potential of quantum computing. Fortunately, quantum error correction[1, 2] provides a promising approach to deal with this issue, through encoding the logical information into certain subspace of the state space of multiple physical qubits. The choice of the subspace corresponds to a quantum error correcting code (QECC).
Quantum low density parity-check (QLDPC) codes[3] are an important class of QECCs. In 2014, Gotteman proved that making use of QLDPC codes, one can realize fault-tolerant quantum computing with constant overhead[4]. Thus, QLDPC codes have attracted a great deal of attention in recent years. So far, there are many construction methods of QLDPC codes have been proposed, such as hypergraph product[5], homological product[6], lifted product[7], balanced product[8] and XYZ product[9]. All the above construction methods generate CSS codes, except XYZ product which generates non-CSS codes.
CSS codes[10] have some advantages. For instance, decoding a CSS code can be transformed into decoding two classical linear codes, which is much easier than decoding a non-CSS code. However, they are not quite good at dealing with biased noise. The stabilizer generators of a CSS code can be generally divided into two parts: one part only contains Pauli $X$ operators (called $X$-type stabilizer) and the other part only contains Pauli $Z$ operators (called $Z$-type stabilizer). $X$-type stabilizers can only detect Pauli $Z$ and $Y$ errors, while $Z$-type stabilizers can detect Pauli $X$ and $Y$ errors. Due to this, under biased noise, some stabilizers of CSS codes cannot provide information of errors, which may result that the error-correcting performance of a CSS code is unsatisfactory under biased noise. Compared to CSS codes, non-CSS codes perform better under biased noise. We take standard surface code[11, 12] and XZZX surface code[13] for an example. For standard surface code, under pure Pauli $Z$ or $X$ noise, only half of stabilizers can provide error-correcting information for decoder, while under pure Pauli $Y$ noise, all stabilizers can provide error-correcting information for decoder. For XZZX surface code, no matter which noise it’s under, the measurement results of all stabilizer can be utilized by decoder. In other words, they all provide decoder with some information about what the actual error looks like. This is one of the important reasons that under pure $X$ and $Z$ noise the error-correcting performance of XZZX surface code is better than that of standard surface code, while under pure $Y$ noise they have similar error correcting performance. Thus, from this perspective, under biased noise, non-CSS codes tend to have better error-correcting performance compared to CSS codes. In the practical hardware system of quantum computing, the probabilities of Pauli $Z$ noise are much higher than those of Pauli $X$ and $Y$ noise, thus considering quantum error-correcting codes with better performance for biased noise is more meaningful.
XYZ product proposed in Ref. [9] constructs a class of non-CSS codes by using three classical codes, which can be seen as a three-fold variant of the hypergraph product or the standard homological product of three classical codes. However, so far there has been no in-depth study of the error-correcting performance of XYZ product codes, and there are several natural questions: whether this construction method can be generalized to higher dimension? If possible, how is the error-correcting performance of these high-dimensional XYZ product codes, especially under biased noise?
The motivation of our work is to answer the problems above. In this paper, we first study the error-correcting performance of three-dimensional (3D) Chamon codes[14, 15], which is a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes with length $n_{1}$, $n_{2}$, $n_{3}$. Exploiting fully decoupled binary belief propagation combined with ordered statistics decoding (FDBP-OSD[16]), when $n_{1}=n_{2}=n_{3}$, the code-capacity threshold of the corresponding Chamon codes is around $14\%$ under depolarizing noise model. For pure Pauli $X$, $Y$ and $Z$ noise model, the corresponding code-capacity threshold of Chamon codes is around $13\%$, $14\%$, $14\%$, respectively. Second, we show that that XYZ product can be generalized to four dimension and propose 4D XYZ product code construction, which can be seen as a variant of standard homological product of two CSS codes (namely, 4D homological product) and constructs a non-CSS code by making use of 4 classical codes or 2 CSS codes. Third, we study the dimension and code distance of 4D XYZ product codes, and show that using the same two CSS codes, the non-CSS code constructed by 4D XYZ product may have higher dimension or code distance than those of the CSS code constructed by 4D homological product. Finally, we explore the error-correcting performance of two special instances of 4D XYZ product. The first one is constructed from two quantum concatenated codes, which are obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$) (here $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd). The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$). We name these two codes 4D XYZ concatenated code and 4D Chamon code, respectively. Exploiting FDBP-OSD, we study their error-correcting performances under the independent single-qubit Pauli noise model. Our simulation results show that under pure Pauli $Z$ noise the code-capacity thresholds of 4D XYZ concatenated codes and 4D Chamon codes are much higher than those of the 4D toric codes[17] and 4D homological concatenated codes (namely, the codes constructed by 4D homological product of two quantum concatenated codes, which are the same two codes used to construct 4D XYZ concatenated code) respectively, while their code-capacity thresholds are close under depolarizing noise. Besides, under $Z$-biased noise model with different bias rate, from the perspective of code-capacity threshold, 4D XYZ concatenated codes and 4D Chamon codes also show good error-correcting performance. Our simulation results means that, using the same two codes, 4D XYZ product can constructed non-CSS codes with better error-correcting performance for biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
This paper is organized as follow. Sect. II introduces some preliminaries, including quantum stabilizer codes, chain complex, hypergraph product codes and 3D XYZ product codes. In Sect. III, we introduces the 4D XYZ product code construction and study its dimension and code distance. In Sect. IV, we explore the error-correcting performance of 3D XYZ Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes. In Sect. V, we conclude our work.
II Preliminaries
In this section, we briefly introduce some preliminaries, including quantum stabilizer codes, chain complex, hypergraph product codes and 3D XYZ product codes.
II.1 Quantum stabilizer codes
Quantum stabilizer codes (QSCs)[18] are an important class of quantum error correcting codes, which can be seen as the analogue of classical linear code in the quantum field.
Giving a $[[n,k]]$ QSC $C$ is equivalent to explicitly giving a set of independent $n$-qubits Pauli operators $S_{1},\cdots,S_{(n-k)}\in\mathcal{P}_{n}$, which commute with each other and are called stabilizer generators, where $\mathcal{P}_{n}$ is the n-qubits Pauli group. This is because the code space $\mathcal{P}_{C}$ of $C$ is the common +1 eigenspace of $S_{1},\cdots,S_{(n-k)}\in\mathcal{P}_{n}$. Formally,
$\mathcal{Q}_{C}=\left\{\ket{\varphi}\in\mathcal{H}_{2}^{\left(\otimes n\right)}:S\ket{\varphi}=\ket{\varphi},\forall S\in\mathcal{S}\right\}$ (1)
where $\mathcal{H}_{2}^{\left(\otimes n\right)}$ is the $n$-qubits Hilbert space and $\mathcal{S}$ is the stabilizer group, which is generated by $S_{1},\cdots,S_{\left(n-k\right)}$, namely $\mathcal{S}=\left\langle S_{1},\cdots,S_{\left(n-k\right)}\right\rangle$.
If the stabilizer generators of a QSC $C$ can be divide into two parts, each of which only contains either Pauli $X$ or Pauli $Z$ operators, it is a CSS code, otherwise it is a non-CSS code. It can be seen that each stabilizer generator of a CSS code can only detect two types of Pauli errors, while that of a non-CSS code can detect all three types of Pauli errors.
The logical operators of a QSC $C$ are the set of operators in $\mathcal{P}_{n}$ which commute with all elements in $\mathcal{S}$ but are not in $\mathcal{S}$. Formally, the logical operators are the elements of $\mathcal{C}\left(\mathcal{S}\right)\backslash\mathcal{S}$, where $\mathcal{C}\left(\mathcal{S}\right)$ is the centralizer of $\mathcal{S}$, which is defined as $\mathcal{C}\left(\mathcal{S}\right)=\left\{P\in\mathcal{P}_{n}:SP=PS,\forall S\in\mathcal{S}\right\}$. For a $\left[\left[n,k,d\right]\right]$ QSC, we can find $k$ pairs of logical operators $\left({\bar{X}}_{j},{\bar{Z}}_{j}\right)_{j=1,\cdots,k}$ such that ${\bar{X}}_{i}{\bar{Z}}_{j}=\left(-1\right)^{\delta_{ij}}{\bar{Z}}_{j}{\bar{X}}_{i}$, where $\delta$ is the Kronecker delta, which means for the same pair of logical operators ${\bar{X}}_{j},{\bar{Z}}_{j}$, they are anti-commute, but they commute with other pairs of logical operators. We can see that $\mathcal{C}\left(\mathcal{S}\right)=\left\langle S_{1},\cdots,S_{\left(n-k\right)},{\bar{X}}_{1},{\bar{Z}}_{1},\cdots,{\bar{X}}_{k},{\bar{Z}}_{k}\right\rangle$. The code distance $d$ is defined as the minimum weight of logical operators, namely,
$d=\mathop{\min}_{L\in\mathcal{C}\left(\mathcal{S}\right)\backslash\mathcal{S}}wt\left(L\right)$ (2)
where the weight of an operator $P\in\mathcal{P}_{n}$ is defined as the number of qubits on which it acts nontrivially, and we use notation $wt\left(\cdot\right)$ to denote it. For instance, $wt\left(I_{1}X_{2}Y_{3}Z_{4}\right)=3$.
II.2 Chain complex
In this section, we introduce the concept of chain complex, which will help to better understand hypergraph product code, 3D XYZ product code and our proposed 4D XYZ product code.
A chain complex $\mathfrak{C}$ with length $L$ is a collection of $L+1$ vector spaces $C_{0},\ C_{1},\cdots,C_{L}$ and $L$ linear maps (which are also called boundary operators) $\partial_{i}:C_{i}\rightarrow C_{i+1}\ \left(0\leq i\leq L-1\right)$, namely,
$\mathfrak{C}=C_{0}\stackrel{{\scriptstyle\partial_{0}}}{{\longrightarrow}}C_{1}\stackrel{{\scriptstyle\partial_{1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\partial_{L-1}}}{{\longrightarrow}}C_{L}$ (3)
which satisfies $\partial_{i}\partial_{i-1}=0$ for all $10\leq i\leq L-$ 1.
If we consider vector space over $\mathbb{F}_{2}$, namely, $C_{i}:=F_{2}^{n_{i}}$, a chain complex $\mathfrak{C}$ with length 2 naturally corresponds to a CSS code $C\left(\mathfrak{C}\right)$, namely,
$C\left(\mathfrak{C}\right)=\mathbb{F}_{2}^{m_{z}}\stackrel{{\scriptstyle H_{z}^{T}}}{{\longrightarrow}}\mathbb{F}_{2}^{N}\stackrel{{\scriptstyle H_{x}}}{{\longrightarrow}}\mathbb{F}_{2}^{m_{x}}$ (4)
where the commutation condition $H_{x}H_{z}^{T}=\mathbf{0}$ is naturally satisfied.
It can be seen that a classical linear code $C=\ker{H}$ corresponds to a chain complex of length 1, namely,
$\mathbb{F}_{2}^{N}\stackrel{{\scriptstyle H}}{{\longrightarrow}}\mathbb{F}_{2}^{m}$ (5)
II.3 Hypergraph product and 3D XYZ product codes
[…]
$\mathfrak{C}=F_{2}^{m_{1}\times n_{2}}\stackrel{{\scriptstyle H_{z}^{T}}}{{\longrightarrow}}F_{2}^{n_{1}\times n_{2}\oplus m_{1}\times m_{2}}\stackrel{{\scriptstyle H_{x}}}{{\longrightarrow}}F_{2}^{n_{1}m_{2}}$ (6)
where $H_{x}=\left(I_{n_{1}}\otimes H_{2},\ H_{1}^{T}\otimes I_{m_{2}}\right)$ and $H_{z}=\left(H_{1}\otimes I_{n_{2}},\ I_{m_{2}}\otimes H_{2}^{T}\right)$. It is easy to verify that $H_{x}H_{z}^{T}=2H_{1}^{T}\otimes H_{2}=\mathbf{0}$. Thus Eq. (6) naturally corresponds to a CSS code $C\left(\mathfrak{C}\right)$. Fig. 1 shows the tensor-product structure corresponding to Eq. (6).
Figure 1: The tensor-product structure of hypergraph product of two classical linear codes $C_{1}=\ker{H_{1}}$ and $C_{2}=\ker{H_{2}}$.
Actually, the tensor-product structure depicted in Fig.1 can not only derive a CSS code, but also derive a non-CSS code. As shown in Fig. 1, the stabilizer generators can be divided into two parts, $S$ and $T$, and each part consists of two classes of qubits, $A$ and $B$. In the case of constructing a CSS code, we choose one between $S$ and $T$ to be $X$-type stabilizers and the other to be $Z$-type stabilizers. If we let stabilizer generators in $S$ and $T$ to be mixed $X$ and $Z$ types, we can obtain a non-CSS code. More precisely, let the nontrivial operators of stabilizers in $S$ acting on qubits in $A$ and $B$ be Pauli $X$ and $Z$ operators, respectively. Similarly, let the nontrivial operators of stabilizers in $T$ acting on qubits in $A$ and $B$ be Pauli $Z$ and $X$ operators, respectively. In this way, we obtain the following stabilizer generator matrix,
$\mathcal{S}=\begin{bmatrix}S\\ T\end{bmatrix}=\begin{bmatrix}X^{\left(H_{1}\otimes I_{n_{2}}\right)},Z^{\left(I_{m_{1}}\otimes H_{2}^{T}\right)}\\ Z^{\left(I_{n_{1}}\otimes H_{2}\right)},X^{\left(H_{1}^{T}\otimes I_{m_{2}}\right)}\end{bmatrix}$ (7)
where the notation $\mathcal{P}=P^{H}\ \left(P\in\left\{X,\ Y,Z\right\}\right)$ denotes a Pauli tensor, which means for any entry of matrix $H$, if it is a 1, $\mathcal{P}$ has a Pauli operator $P$ at the corresponding position, and an identity operator $I$ otherwise. It can be verified any pair of stabilizer generators in $\mathcal{S}$ commutes.
Figure 2: The tensor-product structure of 3D XYZ product of three classical linear code $C_{1}=\ker{H_{1}}$, $C_{2}=\ker{H_{2}}$ and $C_{3}=\ker{H_{3}}$.
[…]
In Fig. 2, $A$, $B$, $C$ and $D$ represent vector spaces which index the qubits and $S$, $T$, $U$ and $V$ represent vector spaces which index the stabilizer generators, namely,
$A\in\mathbb{F}_{2}^{n_{1}\times n_{2}\times n_{3}},\ B\in\mathbb{F}_{2}^{m_{1}\times m_{2}\times n_{3}},$ (8)
$C\in\mathbb{F}_{2}^{m_{1}\times n_{2}\times m_{3}},\ D\in\mathbb{F}_{2}^{n_{1}\times m_{2}\times m_{3}}$
and
$S\in\mathbb{F}_{2}^{m_{1}\times n_{2}\times n_{3}},\ T\in\mathbb{F}_{2}^{n_{1}\times m_{2}\times n_{3}},$ (9)
$U\in\mathbb{F}_{2}^{n_{1}\times n_{2}\times m_{3}},\ V\in\mathbb{F}_{2}^{m_{1}\times m_{2}\times m_{3}}$
Similar with the case of constructing a non-CSS code from the tensor-product structure shown in Fig. 1, the stabilizer generator matrix $\mathcal{S}$ of the corresponding 3D XYZ product code is
$\mathcal{S}=\begin{bmatrix}X^{\left(H_{1}\otimes I_{n_{2}}\otimes I_{n_{3}}\right)}&Y^{\left(I_{m_{1}}\otimes H_{2}^{T}\otimes I_{n_{3}}\right)}&Z^{\left(I_{m_{1}}\otimes I_{n_{2}}\otimes H_{3}^{T}\right)}&I^{\left(m_{1}n_{2}n_{3}\times n_{1}m_{2}m_{3}\right)}\\ Y^{\left(I_{n_{1}}\otimes H_{2}\otimes I_{n_{3}}\right)}&X^{\left(H_{1}^{T}\otimes I_{m_{2}}\otimes I_{n_{3}}\right)}&I^{\left(n_{1}m_{2}n_{3}\times m_{1}n_{2}m_{3}\right)}&Z^{\left(I_{n_{1}}\otimes I_{m_{2}}\otimes H_{3}^{T}\right)}\\ Z^{\left(I_{n_{1}}\otimes I_{n_{2}}\otimes H_{3}\right)}&I^{\left(n_{1}n_{2}m_{3}\times m_{1}m_{2}n_{3}\right)}&X^{\left(H_{1}^{T}\otimes I_{n_{2}}\otimes I_{m_{3}}\right)}&Y^{\left(I_{n_{1}}\otimes H_{2}^{T}\otimes I_{m_{3}}\right)}\\ I^{\left(m_{1}m_{2}m_{3}\times n_{1}n_{2}n_{3}\right)}&Z^{\left(I_{m_{1}}\otimes I_{m_{2}}\otimes H_{3}\right)}&Y^{\left(I_{m_{1}}\otimes H_{2}\otimes I_{m_{3}}\right)}&X^{\left(H_{1}\otimes I_{m_{2}}\otimes I_{m_{3}}\right)}\end{bmatrix}$ (10)
The code length is $N=n_{1}n_{2}n_{3}+m_{1}m_{2}n_{3}+m_{1}n_{2}m_{3}+n_{1}m_{2}m_{3}$.
So far, the error-correcting performance of 3D XYZ product codes has not been studied in depth. In Sect. IV, we study the error-correcting performance of 3D Chamon code, which is a non-CSS variant of the 3D toric code and a special instance of the 3D XYZ product of three repetition codes.
Figure 3: The tensor-product structure of 4D XYZ product of two length-2 chain complexes $\mathfrak{C}_{1}=C_{-1}\stackrel{{\scriptstyle\delta_{-1}}}{{\longrightarrow}}C_{0}\stackrel{{\scriptstyle\delta_{0}}}{{\longrightarrow}}C_{1}$ and $\mathfrak{C}_{2}={\widetilde{C}}_{-1}\stackrel{{\scriptstyle\widetilde{\delta}_{-1}}}{{\longrightarrow}}{\widetilde{C}}_{0}\stackrel{{\scriptstyle\widetilde{\delta}_{0}}}{{\longrightarrow}}{\widetilde{C}}_{1}$.
III Four-dimensional XYZ product codes
In this section, we first introduce the 4D XYZ product code construction. Second, we give out how to compute the dimension of 4D XYZ product codes and compute two special instances of 4D XYZ product codes. Third, we give out the general form of the logical operators of the 4D XYZ product codes, based on which we prove the upper bound of code distance.
III.1 4D XYZ product code construction
4D XYZ product code construction is using either two CSS codes or four classical linear codes to construct a non-CSS code, since one can use four classical linear codes to construct two CSS codes through hypergraph product or concatenation. Formally, the 4D XYZ product code construction is defined as follow.
Definition 1 (4D XYZ product code construction).
Giving two length-2 chain complexes $\mathfrak{C}_{1}=C_{-1}\stackrel{{\scriptstyle\delta_{-1}}}{{\longrightarrow}}C_{0}\stackrel{{\scriptstyle\delta_{0}}}{{\longrightarrow}}C_{1}$ and $\mathfrak{C}_{2}={\widetilde{C}}_{-1}\stackrel{{\scriptstyle\widetilde{\delta}_{-1}}}{{\longrightarrow}}{\widetilde{C}}_{0}\stackrel{{\scriptstyle\widetilde{\delta}_{0}}}{{\longrightarrow}}{\widetilde{C}}_{1}$, which corresponds to two CSS codes $C\left(\mathfrak{C}_{1}\right)$ and $C\left(\mathfrak{C}_{2}\right)$, one can construct the tensor-product structure as shown in Fig. 3, where the qubits are divided into A, B, C, D, E five parts and stabilizer generators are divided into S, T, U, V four parts. The corresponding stabilizer matrix $\mathcal{S}$ is
$\mathcal{S}=\begin{bmatrix}S\\ T\\ U\\ V\end{bmatrix}=\begin{bmatrix}X^{\left(I_{m_{1}}\otimes{\widetilde{\delta}}_{0}^{T}\right)}&Y^{\left(I_{m_{1}}\otimes{\widetilde{\delta}}_{-1}\right)}&Z^{\left(\delta_{-1}^{T}\otimes I_{n_{B}}\right)}&I&I\\ Y^{\left(\delta_{-1}\otimes I_{m_{4}}\right)}&I&X^{\left(I_{n_{A}}\otimes{\widetilde{\delta}}_{0}\right)}&Z^{\left(\delta_{0}^{T}\otimes I_{m_{4}}\right)}&I\\ I&Z^{\left(\delta_{-1}\otimes I_{m_{3}}\right)}&X^{\left(I_{n_{A}}\otimes{\widetilde{\delta}}_{-1}^{T}\right)}&I,&Y^{\left(\delta_{0}^{T}\otimes I_{m_{3}}\right)}\\ I&I&Z^{\left(\delta_{0}\otimes I_{n_{B}}\right)}&Y^{\left(I_{m_{2}}\otimes{\widetilde{\delta}}_{-1}\right)}&X^{\left(I_{m_{2}}\otimes{\widetilde{\delta}}_{-1}\right)}\end{bmatrix}$ (11)
where $m_{1}$, $m_{2}$, $m_{3}$, $m_{4}$, $n_{A}$, $n_{B}$ are the dimensions of $C_{-1}$, $C_{1}$, ${\widetilde{C}}_{-1}$, ${\widetilde{C}}_{1}$, $C_{0}$ and ${\widetilde{C}}_{0}$, respectively. The total number of qubits is
$N=m_{1}m_{4}+m_{1}m_{3}+n_{A}n_{B}+m_{2}m_{4}+m_{2}m_{3}$ (12)
One can see that the tensor-product structure of 4D XYZ product in Fig. 3 is the same as that of 4D homological product, thus it can be seen as a variant of standard 4D homological product of two CSS codes.
Next, we prove that the resulting stabilizer group in Eq. (11) is Abelian.
Corollary 1.
The stabilizer group of the 4D XYZ product code is Abelian.
Proof.
To prove the stabilizer group of the 4D XYZ product code is Abelian is to prove any pair of stabilizers commute. Giving the stabilizer matrix $\mathcal{S}$ of the 4D XYZ product code as shown in Eq. (11), the corresponding parity-check matrix $H$ in symplectic representation is $H=\left(H_{x}\mid H_{z}\right)$, where
$H_{x}=\begin{bmatrix}I_{m_{1}}\otimes{\widetilde{\delta}}_{0}^{T}&I_{m_{1}}\otimes{\widetilde{\delta}}_{-1}&\textbf{0}&\textbf{0}&\textbf{0}\\ \delta_{-1}\otimes I_{m_{4}}&\textbf{0}&I_{n_{A}}\otimes{\widetilde{\delta}}_{0}&\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{0}&I_{n_{A}}\otimes{\widetilde{\delta}}_{-1}^{T}&\textbf{0},&\delta_{0}^{T}\otimes I_{m_{3}}\\ \textbf{0}&\textbf{0}&\textbf{0}&I_{m_{2}}\otimes{\widetilde{\delta}}_{-1}&I_{m_{2}}\otimes{\widetilde{\delta}}_{-1}\end{bmatrix}$ (13)
and
$H_{z}=\begin{bmatrix}\textbf{0}&I_{m_{1}}\otimes{\widetilde{\delta}}_{-1}&\delta_{-1}^{T}\otimes I_{n_{B}}&\textbf{0}&\textbf{0}\\ \delta_{-1}\otimes I_{m_{4}}&\textbf{0}&\textbf{0}&\delta_{0}^{T}\otimes I_{m_{4}}&\textbf{0}\\ \textbf{0}&\delta_{-1}\otimes I_{m_{3}}&\textbf{0}&\textbf{0},&\delta_{0}^{T}\otimes I_{m_{3}}\\ \textbf{0}&\textbf{0}&\delta_{0}\otimes I_{n_{B}}&I_{m_{2}}\otimes{\widetilde{\delta}}_{-1}&\textbf{0}\end{bmatrix}$ (14)
Then we have
$H_{x}H_{z}^{T}+H_{z}H_{x}^{T}$ (15)
$=2\begin{bmatrix}{\widetilde{\delta}}_{-1}{\widetilde{\delta}}_{-1}^{T}&\delta_{-1}^{T}\otimes{\widetilde{\delta}}_{0}^{T}&\delta_{-1}^{T}\otimes{\widetilde{\delta}}_{-1}&\textbf{0}\\ \delta_{-1}\otimes{\widetilde{\delta}}_{0}&\delta_{-1}\delta_{-1}^{T}&\textbf{0}&\delta_{0}^{T}\otimes{\widetilde{\delta}}_{0}^{T}\\ \delta_{-1}\otimes{\widetilde{\delta}}_{-1}^{T}&\textbf{0}&\delta_{0}^{T}\delta_{0}&\delta_{0}^{T}\otimes{\widetilde{\delta}}_{-1}^{T}\\ \textbf{0}&\delta_{0}\otimes{\widetilde{\delta}}_{0}^{T}&\delta_{0}\otimes{\widetilde{\delta}}_{-1}&{\widetilde{\delta}}_{0}^{T}{\widetilde{\delta}}_{0}\end{bmatrix}$
$=\textbf{0}$
which means any pair of stabilizers commute and the proof is completed. ∎
III.2 Dimension of 4D XYZ product code
In this section, we compute the dimension of the 4D XYZ product code. Similar with Ref. [9], this problem is transformed into finding the number of independent stabilizer generators. First, we give a general solution of this problem in Theorem 1. Second, we consider two special instances of 4D XYZ product codes. The first one is constructed from two concatenated code, which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ (here $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and prove that its dimension is 1 in Corollary 2. The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, and prove that its dimension is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$ in Corollary 3. We name these two codes 4D XYZ concatenated code and 4D Chamon code (the reason why we named it this way is explained in Sect. IV.2), respectively.
For convenience, we rewrite Eq. (11) as
$\mathcal{S}=\begin{bmatrix}S\\ T\\ U\\ V\end{bmatrix}=\begin{bmatrix}X^{\left(I_{m_{1}}\otimes H_{x_{2}}^{T}\right)}&Y^{\left(I_{m_{1}}\otimes H_{z_{2}}^{T}\right)}&Z^{\left(H_{z_{1}}\otimes I_{n_{B}}\right)}&I&I\\ Y^{\left(H_{z_{1}}^{T}\otimes I_{m_{4}}\right)}&I&X^{\left(I_{n_{A}}\otimes H_{x_{2}}\right)}&Z^{\left(H_{x_{1}}^{T}\otimes I_{m_{4}}\right)}&I\\ I&Z^{\left(H_{z_{1}}^{T}\otimes I_{m_{3}}\right)}&X^{\left(I_{n_{A}}\otimes H_{z_{2}}\right)}&I,&Y^{\left(H_{x_{1}}^{T}\otimes I_{m_{3}}\right)}\\ I&I&Z^{\left(H_{x1}\otimes I_{n_{B}}\right)}&Y^{\left(I_{m_{2}}\otimes H_{x_{2}}^{T}\right)}&X^{\left(I_{m_{2}}\otimes H_{z_{2}}^{T}\right)}\end{bmatrix}$ (16)
where $H_{x_{1}}=\delta_{0}$ and $H_{z_{1}}=\delta_{-1}^{T}$ are the $X$-type and $Z$-type parity-check matrixes of code $C\left(\mathfrak{C}_{1}\right)$, respectively, and $H_{x_{2}}={\widetilde{\delta}}_{0}$ and $H_{z_{2}}={\widetilde{\delta}}_{-1}^{T}$ are the $X$-type and $Z$-type parity-check matrixes of code $C\left(\mathfrak{C}_{2}\right)$, respectively. Let recall that the dimensions of $H_{x_{1}}$, $H_{z_{1}}$, $H_{x_{2}}$ and $H_{z_{2}}$ are $m_{1}\times n_{A}$, $m_{2}\times n_{A}$, $m_{3}\times n_{B}$ and $m_{4}\times n_{B}$, respectively.
In the rest of paper, we use bold lowercase letters to represent column vectors.
Theorem 1 (The dimension of 4D XYZ product code).
The dimension $k$ of 4D XYZ product code is $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)+k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{s}\\ \textbf{v}\end{bmatrix}$ of
$I_{m_{1}}\otimes H_{x_{2}}\textbf{s}=I_{m_{1}}\otimes H_{z_{2}}\textbf{s}=\textbf{0}$ (17)
$I_{m_{2}}\otimes H_{x_{2}}\textbf{v}=I_{m_{2}}\otimes H_{z_{2}}\textbf{v}=\textbf{0}$
$H_{z_{1}}^{T}\otimes I_{n_{B}}\textbf{s}+H_{x_{1}}^{T}\otimes I_{n_{B}}\textbf{v}=\textbf{0}$
and $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{t}\\ \textbf{u}\end{bmatrix}$ of
$H_{z_{1}}\otimes I_{m_{4}}\textbf{t}=H_{x_{1}}\otimes I_{m_{4}}\textbf{t}=\textbf{0}$ (18)
$H_{z_{1}}\otimes I_{m_{3}}\textbf{u}=H_{z_{1}}\otimes I_{m_{3}}\textbf{u}=\textbf{0}$
$H_{z_{1}}\otimes I_{m_{3}}\textbf{u}=H_{z_{1}}\otimes I_{m_{3}}\textbf{u}=\textbf{0}$
Proof.
The total number $m$ of stabilizer generators in Eq. (16) is $m=m_{1}n_{B}+m_{2}n_{B}+m_{3}n_{A}+m_{4}n_{A}$. Thus, according to Eq. (12), the 4D XYZ product code at least encodes $N-m=\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)$ logical qubits. However, we should notice that stabilizer generators in Eq. (16) may not be independent with each other. Thus, we should find out the number of independent stabilizer generators in Eq. (16).
Observing Eq. (16), one can see that the stabilizer generators are divided into $S$, $T$, $U$, $V$ four parts, and any stabilizer in $S$ or $V$ cannot be generated by the product of some stabilizers in $T$ and $U$. Thus, we first consider the independence of the stabilizers in $S$ and $V$, and for the stabilizers in $T$ and $U$ the analysis is the same.
Considering stabilizers in $S$ and $V$ simultaneously, one can see that the product of some stabilizers in $S$ might be equal to the product of some stabilizers in $V$, namely,
$\left[I_{m_{1}}\otimes H_{x_{2}}^{T},\ I_{m_{1}}\otimes H_{z_{2}}^{T},\ H_{z_{1}}\otimes I_{n_{B}},\ \textbf{0},\ \textbf{0}\right]^{T}\textbf{\emph{s}}$ (19)
$=\left[\textbf{0},\ \textbf{0},\ H_{x_{1}}\otimes I_{n_{B}},\ I_{m_{2}}\otimes H_{x_{2}}^{T},\ I_{m_{2}}\otimes H_{z_{2}}^{T}\right]^{T}\textbf{\emph{v}}$
Then we have Eq. (17).
Our goal is to compute $k_{SV}$ and $k_{TU}$. Observing that equation systems (17) and (18) are similar, thus we only need to compute $k_{SV}$, and $k_{TU}$ will follow a similar fashion.
It can be seen that the solution of $I_{m_{1}}\otimes H_{x_{2}}\textbf{\emph{s}}=I_{m_{1}}\otimes H_{z_{2}}\textbf{\emph{v}}=\mathbf{0}$ is $\textbf{\emph{s}}\in\mathcal{C}_{m_{1}}\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ and the solution of $I_{m_{2}}\otimes H_{x_{2}}\textbf{\emph{v}}=I_{m_{2}}\otimes H_{z_{2}}\textbf{\emph{v}}=\mathbf{0}$ is $\textbf{\emph{v}}\in\mathcal{C}_{m_{2}}\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$. Similarly, the solution of $H_{z_{1}}^{T}\otimes I_{n_{B}}\textbf{\emph{s}}=H_{x_{1}}^{T}\otimes I_{n_{B}}\textbf{\emph{v}}$ is $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\mathcal{C}_{n_{B}}$. Thus, the solution of equation system (17) is $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, and the dimension of $\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}$ of equation system (17).
Let $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, according to the theory of linear algebra, the number of independent stabilizers $r_{SV}$ is
$r_{SV}=m_{1}n_{B}+m_{2}n_{B}-k_{SV}$ (20)
Similarly, let $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{\emph{t}}\\ \textbf{\emph{u}}\end{bmatrix}$, the number of independent stabilizers $r_{TU}$ is
$r_{TU}=m_{3}n_{A}+m_{4}n_{A}-k_{TU}$ (21)
Thus, the total number of independent stabilizers is $r_{SV}+r_{TU}$, and the dimension of 4D XYZ product code is
$N-\left(m_{1}n_{B}+m_{2}n_{B}-k_{SV}\right)-\left(m_{3}n_{A}+m_{4}n_{A}-k_{TU}\right)$ (22)
$=\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)+k_{SV}+k_{TU}$
Next, we consider 4D XYZ product of two concatenated codes obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ ($n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and compute it dimension. Let the repetition code $C_{1}$ with length $n_{1}$ be the outer code and the code $C_{2}$ with length $n_{2}$ be the inner code, whose parity-check matrices are $H_{1}$ with dimension $\left(n_{1}-1\right)\times n_{1}$ and $H_{2}$ with dimension $\left(n_{2}-1\right)\times n_{2}$ respectively, the corresponding $X$-type and $Z$-type parity-check matrices $H_{x}$ and $H_{z}$ of the concatenated code obtained from $C_{1}$ and $C_{2}$ have the following form:
$H_{x}=\begin{bmatrix}\overbrace{1,\cdots,1}^{n_{2}},&\overbrace{1,\cdots,1}^{n_{2}},&\overbrace{0,\cdots,0}^{(n_{1}-2)n_{2}}\\ \vdots&\ddots&\vdots\\ \underbrace{0,\cdots,0}_{(n_{1}-2)n_{2}},&\underbrace{1,\cdots,1}_{n_{2}},&\underbrace{1,\cdots,1}_{n_{2}}\end{bmatrix}$ (23)
$H_{z}=\begin{bmatrix}H_{1}&\cdots&\textbf{0}\\ \vdots&\ddots&\vdots\\ \textbf{0}&\cdots&H_{1}\\ \end{bmatrix}$ (24)
Corollary 2 (The dimension of 4D XYZ concatenated code).
The dimension $k$ of the code, which is 4D XYZ product of two concatenated codes obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ (where $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), is 1.
Proof.
According to Theorem 1 and $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)=1$, the dimension $k$ of 4D XYZ concatenated code is $1+k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$.
Observing that $\dim\left(\ker\left(\begin{bmatrix}H_{x}\\ H_{z}\end{bmatrix}\right)\right)$ can be seen as the number of independent solutions $\hat{\textbf{\emph{e}}}=\left(\textbf{\emph{e}},\textbf{\emph{e}}\right)^{T}$ of
$\begin{pmatrix}H_{x}&\textbf{0}\\ \textbf{0}&H_{z}\end{pmatrix}\left(\textbf{\emph{e}},\textbf{\emph{e}}\right)^{T}=\begin{pmatrix}\textbf{0}\\ \textbf{0}\end{pmatrix}$ (25)
Here, the solution $\hat{\textbf{\emph{e}}}$ can be interpreted as the symplectic representation of undetectable $Y$-type errors, which include $Y$-type stabilizers and logical operators of the corresponding code.
For a concatenated code $C$ obtained from a pair of repetition codes, whose $X$-type and $Z$-type parity-check matrices are $H_{x}$ and $H_{z}$ respectively, it is easy to prove that $\dim\left(\ker\left(\begin{bmatrix}H_{x}\\ H_{z}\end{bmatrix}\right)\right)=1$ and $\dim\left(\ker\left(\left[H_{z}^{T},H_{x}^{T}\right]\right)\right)=n-1-dim\left(row\left(\begin{bmatrix}H_{x}\\ H_{z}\end{bmatrix}\right)\right)=n-1-(n-1)=0$. The reason is that, for code $C$ the undetectable $Y$-type error only contain one $Y$-type logical operator whose weight is code length. Thus, the dimension $k$ of 4D XYZ concatenated code is 1. ∎
Next, we prove Lemma 1 which will be used in the proof of Corollary 3.
Lemma 1 (The number of independent undetectable $Y$-type errors of 2D toric code).
The number of independent $Y$-type stabilizers and logical operators of 2D toric code is $2\gcd{\left(j,k\right)}$, where $j$ and $k$ are dimensions of the toric code lattice.
Proof.
Without loss of generality, we assume that $j\leq k$, and the proof consists of two parts:
1. Proving that the number of independent $Y$-type logical operators of 2D toric code is 2.
2. Proving that the number of independent $Y$-type stabilizers of 2D toric code is $2\left[gcd\left(j,k\right)-1\right]$.
Since 2D toric code encodes 2 logical operators, there are 2 independent $Y$-type logical operators at most. When the toric code lattice is a square, these two independent $Y$-type logical operators are full diagonal of Pauli $Y$ operators from the upper right to the lower left and from the upper left to the lower right, respectively. In Fig. 4, we give an example of $Y$-type logical operators of 2D toric code with lattice size $3\times 3$.
Figure 4: $Y$-type logical operators of 2D toric code with lattice size $3\times 3$.
Any $Y$-type stabilizer must be generated by the product of some $X$-type and $Z$-type stabilizers both of which act on the same set of qubits. According to the topology of toric code, this can only be realized by selecting plaquettes ($Z$-type stabilizers) and vertices ($X$-type stabilizers) along the diagonal of $j\times j$ square lattices periodically. Fig. 5 gives an example of $Y$-type stabilizer of toric code with dimension $j=4,\ k=6$.
Figure 5: A $Y$-type stabilizer of toric code with lattice size $j=4,\ k=6$. The red solid line edges are qubits on which the $Y$-type stabilizer acts. Each plaquette filled with blue-grey is the selected $Z$-type stabilizer, and its top right and bottom left vertices are the selected $X$-type stabilizers.
One can see that when $j$ and $k$ are coprime, namely, $\gcd\left(j,k\right)=1$, we cannot find a selection method of $X$-type and $Z$-type stabilizers to generate a $Y$-type stabilizer, thus the number of $Y$-type stabilizers is zero. While in the case of $\gcd\left(j,k\right)=g\neq 1$, using the selection method as shown in Fig. 5, one can find $2\left(g-1\right)$ independent $Y$-type stabilizers by shifting to left or right, where the factor 2 comes from the left-right symmetry of toric code.
To sum up, the total number of independent $Y$-type stabilizers and logical operators of 2D toric code is $2\gcd{\left(j,k\right)}$, and the proof is completed. ∎
Corollary 3 (The dimension of the 4D Chamon code).
The dimension $k$ of the 4D Chamon code, which is 4D XYZ product of two hypergraph product codes that are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, is
$k=8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$ (26)
Proof.
The hypergraph product of a pair of repetition codes is a toric code[5]. Thus, given two toric codes which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, whose parity-check matrixes are $H_{1}$, $H_{2}$, $H_{3}$ and $H_{4}$, respectively, by Definition 1, the corresponding stabilizer matrix s of the 4D XYZ product code is as shown in Eq. (16), where $H_{x_{1}}=\left[I_{n_{1}}\otimes H_{2},H_{1}^{T}\otimes I_{n_{2}}\right]$ with dimension of $m_{1}\times n_{A}$ and $H_{z_{1}}=\left[H_{1}\otimes I_{n_{2}},\ I_{n_{1}}\otimes H_{2}^{T}\right]$ with dimension of $m_{2}\times n_{A}$ are the $X$-type and $Z$-type parity-check matrixes of toric code $C_{1}$, respectively, and $H_{x_{2}}=\left[I_{n_{3}}\otimes H_{4},H_{3}^{T}\otimes I_{n_{4}}\right]$ with dimension of $m_{3}\times n_{B}$ and $H_{z_{2}}=\left[H_{3}\otimes I_{n_{4}},\ I_{n_{3}}\otimes H_{4}^{T}\right]$ with dimension of $m_{4}\times n_{B}$ are the $X$-type and $Z$-type parity-check matrixes of toric code $C_{2}$, respectively. Let recall that $n_{A}=2n_{1}n_{2}=m_{1}+m_{2}$ and $n_{B}=2n_{3}n_{4}=m_{3}+m_{4}$ is the code length of $C_{1}$ and $C_{2}$, respectively.
According to Theorem 1 and $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)=0$, the dimension $k$ of the 4D Chamon code is $k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$.
Similar with the proof in Corollary 2, $\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is the number of undetectable $Y$-type errors, which include $Y$-type stabilizers and logical operators of the toric code $C_{2}$. According to Lemma 1, the number of independent solutions $\hat{\textbf{\emph{e}}}$ is $2\gcd(n_{3},n_{4})$.
For $\dim{\left(\ker{\left(\left[H_{z1}^{T},\ \ H_{x1}^{T}\right]\right)}\right)}$, it is equal to $n_{A}-\dim{\left(Im\left(\left[H_{z_{1}}^{T},\ \ H_{x_{1}}^{T}\right]\right)\right)}=n_{A}-row\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)=\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)=2\gcd{\left(n_{1},n_{2}\right)}$, where $Im\left(\left[H_{z_{1}}^{T},\ \ H_{x_{1}}^{T}\right]\right)$ is the image space of $\left[H_{z_{1}}^{T},\ \ H_{x_{1}}^{T}\right]$ and $row\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)$ is the row space of $\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}$.
Observing that the dimension $k_{SV}$ of $\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is $\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, thus $k_{SV}=4\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$. Similarly, the dimension $k_{TU}$ of $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$ is also $4\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$. Thus, the dimension $k$ of the 4D Chamon code is $k=k_{SV}+k_{TU}=8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, which completes the proof. ∎
For 4D toric code which is 4D homological product of two 2D toric codes, no matter what values $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ take, its dimension is always 6. While the dimension of 4D Chamon code, $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, is higher than that of 4D toric code, which means using the same two CSS codes, 4D XYZ product may construct codes with higher dimension than 4D homological product.
III.3 Logical operators and code distance of 4D XYZ product code
In this section, we study the logical operators and the code distance of 4D XYZ product code. First, we give out the general form of logical operators of the 4D XYZ product code in Theorem 2, then we compute their minimum weight in corollary 4.
Here we define a permutation notation $\pi\left(\cdot\right)$ as
$\pi\left(\begin{bmatrix}x_{1}\textbf{\emph{a}}\\ \vdots\\ x_{m_{1}}\textbf{\emph{a}}\\ x_{1}\textbf{\emph{b}}\\ \vdots\\ x_{m_{1}}\textbf{\emph{b}}\\ y_{1}\textbf{\emph{a}}\\ \vdots\\ y_{m_{2}}\textbf{\emph{a}}\\ y_{1}\textbf{\emph{b}}\\ \vdots\\ y_{m_{2}}\textbf{\emph{b}}\end{bmatrix}\right)=\begin{bmatrix}x_{1}\textbf{\emph{a}}\\ x_{1}\textbf{\emph{b}}\\ \vdots\\ x_{m_{1}}\textbf{\emph{a}}\\ x_{m_{1}}\textbf{\emph{b}}\\ y_{1}\textbf{\emph{a}}\\ y_{1}\textbf{\emph{b}}\\ \vdots\\ y_{m_{2}}\textbf{\emph{a}}\\ y_{m_{2}}\textbf{\emph{b}}\end{bmatrix}$ (27)
which will be used in the following proof.
Theorem 2 (The logical operators of 4D XYZ product code).
Giving two CSS codes $C_{1}$ and $C_{2}$, whose $X$-type and $Z$-type parity-check matrixes are $H_{x_{1}}$, $H_{z_{1}}$ and $H_{x_{2}}$, $H_{z_{2}}$, respectively, the logical operators of the corresponding 4D XYZ product code are divided into two types. The first type of logical operators are
$X_{L}=\left[I\ \ \ \ \ I\ \ \ \ X\left(\textbf{r}^{T}\right)\ \ \ I\ \ \ \ \ I\right]$ (28)
$Z_{L}=\left[I\ \ \ \ \ I\ \ \ \ Z\left(\textbf{w}^{T}\right)\ \ \ I\ \ \ \ \ I\right]$
where
$\textbf{r}\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\left(\mathcal{C}_{n_{B}}\backslash Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$ (29)
$\textbf{w}\in\left(\mathcal{C}_{n_{A}}\backslash Im\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$
The second type of logical operators are
$\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\\ Z_{L_{2}}\\ Z_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{a}_{1})\ &I\ &I\ &Y(\textbf{b}_{1})\ &I\\ I\ &Y(\textbf{a}_{2})\ &I\ &I\ &X(\textbf{b}_{2})\\ Y(\textbf{c}_{1})\ &Z(\textbf{d}_{1})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{c}_{2})\ &Y(\textbf{d}_{2})\end{bmatrix}$ (30)
where
${\pi\left(\textbf{a}_{1}^{T},\ \textbf{a}_{2}^{T},\textbf{b}_{1}^{T},\textbf{b}_{2}^{T}\right)}^{T}$ (31)
$\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$
${\pi\left(\textbf{c}_{1}^{T},\ \textbf{c}_{2}^{T},\textbf{d}_{1}^{T},\textbf{d}_{2}^{T}\right)}^{T}$
$\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$
Proof.
Our goal is to find out all independent logical operators, and our method is divided into three steps. First, finding out some operators that commutes with all stabilizer generators in Eq. (16). Second, excluding those operators which are stabilizers from the operators that we find in the first step, and the remaining operators are logical operators. Finally, we prove that the number of independent logical operator pairs that we find in the second step is equal to the dimension of the corresponding 4D XYZ product code that we prove in Theorem 1.
First, we consider the operators with the following form, which commute with all stabilizer generators,
$X_{G_{1}}=\left[I\ \ \ \ \ I\ \ \ \ X\left({\textbf{\emph{r}}^{\prime}}^{\textbf{T}}\right)\ \ \ I\ \ \ \ \ I\right]$ (32)
$Z_{G_{1}}=\left[I\ \ \ \ \ I\ \ \ \ X\left({\textbf{\emph{w}}^{\prime}}^{\textbf{T}}\right)\ \ \ I\ \ \ \ \ I\right]$
where ${\textbf{\emph{r}}^{\prime}}^{T}\in\ker\left(\begin{bmatrix}H_{x_{1}}\otimes I_{n_{B}}\\ H_{z_{1}}\otimes I_{n_{B}}\end{bmatrix}\right)=\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\mathcal{C}_{n_{B}}$ and ${\textbf{\emph{w}}^{\prime}}^{T}\in\ker\left(\begin{bmatrix}I_{n_{A}}\otimes H_{x_{2}}\\ I_{n_{A}}\otimes H_{z_{2}}\end{bmatrix}\right)=\mathcal{C}_{n_{A}}\otimes\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)$.
Let ${\textbf{\emph{r}}^{\prime\prime}}^{T}\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes Im\left(H_{x_{2}}^{T},H_{z_{2}}^{T}\right)\subset\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\mathcal{C}_{n_{B}}$, next we prove that $X_{G_{1}^{\prime}}=\left[I\ I\ X\left({\textbf{\emph{r}}^{\prime\prime}}^{T}\right)\ I\ I\right]$ is a stabilizer which can only be generated by the product of some stabilizer generators in $T$ and $U$ and thus should be excluded. Considering an operator $P_{1}$ which is the product of some stabilizer generators in $T$ and $U$, it must have the following form
$P_{1}$ $=\left[\textbf{\emph{t}}^{T},\textbf{\emph{u}}^{T}\right]\begin{bmatrix}Y^{\left(H_{z_{1}}^{T}\otimes I_{m_{4}}\right)}&I&X^{\left(I_{n_{A}}\otimes H_{x_{2}}\right)}&Z^{\left(H_{x_{1}}^{T}\otimes I_{m_{4}}\right)}&I\\ I&Z^{\left(H_{z_{1}}^{T}\otimes I_{m_{3}}\right)}&X^{\left(I_{n_{A}}\otimes H_{z_{2}}\right)}&I,&Y^{\left(H_{x_{1}}^{T}\otimes I_{m_{3}}\right)}\end{bmatrix}$ (33)
$=\left[Y^{\textbf{\emph{t}}^{T}\left(H_{z_{1}}^{T}\otimes I_{m_{4}}\right)}\ \ \ \ Z^{\textbf{\emph{u}}^{T}\left(H_{z_{1}}^{T}\otimes I_{m_{3}}\right)}\ \ \ \ X^{\textbf{\emph{t}}^{T}\left(I_{n_{A}}\otimes H_{x_{2}}\right)+\textbf{\emph{u}}^{T}\left(I_{n_{A}}\otimes H_{z_{2}}\right)}\ \ \ \ Z^{\textbf{\emph{t}}^{T}\left(H_{x_{1}}^{T}\otimes I_{m_{4}}\right)}\ \ \ \ Y^{\textbf{\emph{t}}^{T}\left(H_{x_{1}}^{T}\otimes I_{m_{3}}\right)}\right]$
If the form of $P_{1}$ is the same with that of $X_{G_{1}^{\prime}}$, we must have
$H_{z_{1}}\otimes I_{m_{4}}\textbf{\emph{t}}=H_{x1}\otimes I_{m_{4}}\textbf{\emph{t}}=\mathbf{0}$ (34)
$H_{z1}\otimes I_{m_{3}}\textbf{\emph{u}}=H_{x1}\otimes I_{m_{3}}\textbf{\emph{u}}=\mathbf{0}$
[…]
$\left(I_{n_{A}}\otimes H_{x_{2}}^{T}\right)\left(\textbf{\emph{x}}\otimes\textbf{\emph{i}}\right)+\left(I_{n_{A}}\otimes H_{z_{2}}^{T}\right)\left(\textbf{\emph{y}}\otimes\textbf{\emph{j}}\right)$ (35)
$=\textbf{\emph{x}}\otimes H_{x_{2}}^{T}\textbf{\emph{i}}+\textbf{\emph{y}}\otimes H_{z_{2}}^{T}\textbf{\emph{j}}$
$\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes Im\left(H_{x_{2}}^{T},H_{z_{2}}^{T}\right)$
Thus, for vector $\textbf{\emph{r}}\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\left(\mathcal{C}_{n_{B}}\backslash Im\left(H_{x_{2}}^{T},H_{z_{2}}^{T}\right)\right)$, the corresponding operator $X_{L}=\left(I\ I\ X\left(\textbf{\emph{r}}^{T}\right)\ I\ I\right)$ is a logical operator.
Similarly, for vector $\textbf{\emph{w}}^{\prime\prime}\in Im\left(\left[H_{x1}^{T},H_{z1}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, the corresponding $Z_{G_{1}^{\prime}}=\left(I\ I\ Z\left({\textbf{\emph{w}}^{\prime\prime}}^{T}\right)\ I\ I\right)$ is a stabilizer which can only be generated by the product of some stabilizer generators in $S$ and $V$ and thus should be excluded.
To sum up, for vectors $\textbf{\emph{r}}\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\left(\mathcal{C}_{n_{B}}\backslash Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$ and $\textbf{\emph{w}}\in\left(\mathcal{C}_{n_{A}}\backslash Im\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, the corresponding operators $X_{L}=\left(I\ I\ X\left(\textbf{\emph{r}}^{T}\right)\ I\ I\right)$ and $Z_{L}=\left(I\ I\ Z\left(\textbf{\emph{w}}^{T}\right)\ I\ I\right)$ must be logical operators.
Notice that the dimension of vector space $\mathcal{C}_{n_{B}}\backslash Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)$ is $n_{B}-\dim\left(Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)=n_{B}-\dim\left(row\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)=\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$. Similarly, the dimension of vector space $\mathcal{C}_{n_{A}}\backslash Im\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)$ is equal to $n_{A}-\dim\left(row\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)$. Thus, the numbers of independent logical operators $X_{L}$ and $Z_{L}$ are both $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$. Moreover, for any vector $\textbf{\emph{r}}\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\left(\mathcal{C}_{n_{B}}\backslash Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$, we can find a vector $\textbf{\emph{w}}\in\left(\mathcal{C}_{n_{A}}\backslash Im\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, such that $\textbf{\emph{r}}\cdot\textbf{\emph{w}}^{T}=1$, which means the corresponding $X_{L}$ and $Z_{L}$ anti-commute. Thus, there are $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ pairs of the first type logical operators.
Second, we consider operators with the following form, which also commute with all stabilizer generators,
$\begin{bmatrix}X_{G_{2}}\\ X_{G_{3}}\\ Z_{G_{2}}\\ Z_{G_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1}^{\prime})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1}^{\prime})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2}^{\prime})\ &I\ &I\ &X(\textbf{\emph{b}}_{2}^{\prime})\\ Y(\textbf{\emph{c}}_{1}^{\prime})\ &Z(\textbf{\emph{d}}_{1}^{\prime})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{\emph{c}}_{2}^{\prime})\ &Y(\textbf{\emph{d}}_{2}^{\prime})\end{bmatrix}$ (36)
where $\left({\textbf{\emph{a}}_{1}^{\prime}}^{T},\ {\textbf{\emph{b}}_{1}^{\prime}}^{T}\right)^{T}\in\ker{\left(\left[H_{z_{1}}^{T}\otimes I_{m_{4}},H_{x_{1}}^{T}\otimes I_{m_{4}}\right]\right)}=\ker{\left(\left[H_{z1}^{T},H_{x1}^{T}\right]\right)}\otimes\mathcal{C}_{m_{4}}$ and $\left({\textbf{\emph{a}}_{2}^{\prime}}^{T},\ {\textbf{\emph{b}}_{2}^{\prime}}^{T}\right)^{T}\in\ker{\left(\left[H_{z_{1}}^{T}\otimes I_{m_{3}},H_{x_{1}}^{T}\otimes I_{m_{3}}\right]\right)}=\ker{\left(\left[H_{z1}^{T},H_{x1}^{T}\right]\right)}\otimes\mathcal{C}_{m_{3}}$.
Let ${\pi\left({\textbf{\emph{a}}_{1}^{\prime\prime}}^{T},\ {\textbf{\emph{a}}_{2}^{\prime\prime}}^{T},{\textbf{\emph{b}}_{1}^{\prime\prime}}^{T},{\textbf{\emph{b}}_{2}^{\prime\prime}}^{T}\right)}^{T}\in\ker{\left(\left[H_{z1}^{T},H_{x1}^{T}\right]\right)}\otimes Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ and $\begin{bmatrix}X_{G_{2}}^{\prime\prime}\\ X_{G_{3}}^{\prime\prime}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1}^{\prime\prime})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1}^{\prime\prime})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2}^{\prime\prime})\ &I\ &I\ &X(\textbf{\emph{b}}_{2}^{\prime\prime})\end{bmatrix}$. Next, we prove that $X_{G_{2}^{\prime\prime}}X_{G_{3}^{\prime\prime}}=\left[X\left({\textbf{\emph{a}}_{1}^{\prime\prime}}^{T}\right)\ Y\left({\textbf{\emph{a}}_{2}^{\prime\prime}}^{T}\right)\ I\ Y\left({\textbf{\emph{b}}_{1}^{\prime\prime}}^{T}\right)\ X\left({\textbf{\emph{b}}_{2}^{\prime\prime}}^{T}\right)\right]$ is a stabilizer, which can only be generated by the product of some stabilizer generators in $S$ and $V$, thus should be excluded. Considering an operator $P_{2}$ which is the product of some stabilizer generators in $S$ and $V$, it must have the following form
$P_{2}$ $=\left[\textbf{\emph{s}}^{T},\textbf{\emph{t}}^{T}\right]\begin{bmatrix}X^{\left(I_{m_{1}}\otimes H_{x_{2}}^{T}\right)}&Y^{\left(I_{m_{1}}\otimes H_{z_{2}}^{T}\right)}&Z^{\left(H_{z_{1}}\otimes I_{n_{B}}\right)}&I&I\\ I&I&Z^{\left(H_{x1}\otimes I_{n_{B}}\right)}&Y^{\left(I_{m_{2}}\otimes H_{x_{2}}^{T}\right)}&X^{\left(I_{m_{2}}\otimes H_{z_{2}}^{T}\right)}\end{bmatrix}$ (37)
$=\left[X^{\textbf{\emph{s}}^{T}\left(I_{m_{1}}\otimes H_{x_{2}}^{T}\right)}\ \ \ \ \ Y^{\textbf{\emph{s}}^{T}\left(I_{m_{1}}\otimes H_{z_{2}}^{T}\right)}\ \ \ \ Z^{\textbf{\emph{s}}^{T}\left(H_{z_{1}}\otimes I_{n_{B}}\right)+\textbf{\emph{v}}^{T}\left(H_{x1}\otimes I_{n_{B}}\right)}\ \ \ \ Y^{\textbf{\emph{v}}^{T}\left(I_{m_{2}}\otimes H_{x_{2}}^{T}\right)}\ \ \ \ X^{\textbf{\emph{v}}^{T}\left(I_{m_{2}}\otimes H_{z_{2}}^{T}\right)}\right]$
If the form of $P_{2}$ is the same with $X_{G_{2}^{\prime\prime}}X_{G_{3}^{\prime\prime}}$, we must have
$H_{z1}^{T}\otimes I_{n_{B}}\textbf{\emph{s}}+H_{x1}^{T}\otimes I_{n_{B}}\textbf{\emph{v}}=\mathbf{0}$ (38)
Thus, we have $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z1}^{T}\otimes I_{n_{B}},H_{x1}^{T}\otimes I_{n_{B}}\right]\right)=\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\mathcal{C}_{n_{B}}$. Let $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}=\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}\otimes\textbf{\emph{i}}$, where $\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}=\left[x_{1},\cdots,x_{m_{1}},y_{1},\cdots,y_{m_{y}}\right]^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}$ and $\textbf{\emph{i}}\in\mathcal{C}_{n_{B}}$, we have
$\begin{bmatrix}I_{m_{1}}\otimes H_{x_{2}}&\textbf{0}\\ I_{m_{1}}\otimes H_{z_{2}}&\textbf{0}\\ \textbf{0}&I_{m_{2}}\otimes H_{x_{2}}\\ \textbf{0}&I_{m_{2}}\otimes H_{z_{2}}\end{bmatrix}\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}\otimes\textbf{\emph{i}}=\begin{bmatrix}\textbf{\emph{x}}\otimes H_{x_{2}}\textbf{\emph{i}}\\ \textbf{\emph{x}}\otimes H_{z_{2}}\textbf{\emph{i}}\\ \textbf{\emph{y}}\otimes H_{x_{2}}\textbf{\emph{i}}\\ \textbf{\emph{y}}\otimes H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}$ (39)
$=\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}$
It can be seen that
$\pi\left(\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}\right)=\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}$ (40)
$\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}$
which means $X_{G_{2}^{\prime\prime}}X_{G_{3}^{\prime\prime}}$ is a stabilizer. Thus, for vectors ${\pi\left(\textbf{\emph{a}}_{1}^{T},\textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, the corresponding operators $\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2})\ &I\ &I\ &X(\textbf{\emph{b}}_{2})\end{bmatrix}$ are logical operators.
Similarly, for vector $\left({\textbf{\emph{c}}_{1}^{\prime}}^{T},{\textbf{\emph{d}}_{1}^{\prime}}^{T}\right)^{T}\in\ker{\left(I_{m_{1}}\otimes H_{x_{2}}^{T},I_{m_{1}}\otimes H_{z2}^{T}\right)}$, $\left({\textbf{\emph{c}}_{2}^{\prime}}^{T},\ {\textbf{\emph{d}}_{2}^{\prime}}^{T}\right)^{T}\in\ker{\left(I_{m_{2}}\otimes H_{x2}^{T},I_{m_{2}}\otimes H_{z2}^{T}\right)}$ and $\begin{bmatrix}Z_{G_{2}}^{\prime\prime}\\ Z_{G_{3}}^{\prime\prime}\end{bmatrix}=\begin{bmatrix}Y(\textbf{\emph{c}}_{1}^{\prime\prime})\ &Z(\textbf{\emph{d}}_{1}^{\prime\prime})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{\emph{c}}_{2}^{\prime\prime})\ &Y(\textbf{\emph{d}}_{2}^{\prime\prime})\end{bmatrix}$, the operator $Z_{G_{2}^{\prime\prime}}Z_{G_{3}^{\prime\prime}}=\left[Y\left({\textbf{\emph{c}}_{1}^{\prime\prime}}^{T}\right)\ \ Z\left({\textbf{\emph{c}}_{2}^{\prime\prime}}^{T}\right)\ \ I\ \ Z\left({\textbf{\emph{d}}_{1}^{\prime\prime}}^{T}\right)\ \ Y\left({\textbf{\emph{d}}_{2}^{\prime\prime}}^{T}\right)\right]$ is a stabilizer which can only be generated by the product of some stabilizer generators in $T$ and $U$ and thus should be excluded.
To sum up, for vectors ${\pi\left(\textbf{\emph{a}}_{1}^{T},\ \textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and ${\pi\left(\textbf{\emph{c}}_{1}^{T},\ \textbf{\emph{c}}_{2}^{T},\textbf{\emph{d}}_{1}^{T},\textbf{\emph{d}}_{2}^{T}\right)}^{T}\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$, the corresponding $\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2})\ &I\ &I\ &X(\textbf{\emph{b}}_{2})\end{bmatrix}$ and $\begin{bmatrix}Z_{L_{2}}\\ Z_{L_{3}}\end{bmatrix}=\begin{bmatrix}Y(\textbf{\emph{c}}_{1})\ &Z(\textbf{\emph{d}}_{1})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{\emph{c}}_{2})\ &Y(\textbf{\emph{d}}_{2})\end{bmatrix}$ must be logical operators.
Notice that the dimension of vector space $\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}$ is $m_{3}+m_{4}-\dim\left(Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)=m_{3}+m_{4}-dim\left(row\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)=\dim\left(\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$. Similarly, the dimension of vector space $\mathcal{C}_{m_{1}+m_{1}}\backslash Im\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}$ is $m_{1}+m_{2}-\dim\left(Im\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)=m_{1}+m_{2}-dim\left(row\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)=\dim\left(\ker\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)$. Thus, the total number of independent logical operators $X_{L_{2}}$ and $X_{L_{3}}$ and that of $Z_{L_{2}}$ and $Z_{L_{3}}$ are both $\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$. Moreover, for any vector $\textbf{\emph{r}}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, we can find a vector $\textbf{\emph{w}}\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$, such that $\textbf{\emph{r}}\cdot\textbf{\emph{w}}^{T}=1$, which means the corresponding $X_{L}$ and $Z_{L}$ anti-commute. Thus, there are $\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$ pairs of the second type logical operators.
The total number of the first and the second types of logical operators is $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)+\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$. Now, we prove that this number is equal to the dimension of the corresponding 4D XYZ product code as we prove in Theorem 1.
Let $a=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)$ and $b=\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, we have $\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)=m_{1}+m_{2}-\dim\left(row{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)=m_{1}+m_{2}-\left(n_{A}-\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\right)=m_{1}+m_{2}-(n_{A}-a)$. Similarly, $\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)=m_{3}+m_{4}-\dim\left(row{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)=m_{3}+m_{4}-\left(n_{B}-\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)\right)=m_{3}+m_{4}-(n_{B}-b)$. Thus,
$\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ (41)
$+\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$
$=ab+\left[m_{1}+m_{2}-\left(n_{A}-a\right)\right]\left[m_{3}+m_{4}-\left(n_{B}-b\right)\right]$
$=a\left[b-\left(n_{B}-m_{3}-m_{4}\right)\right]+b\left[a-\left(n_{A}-m_{1}-m_{2}\right)\right]$
$+\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)$
which is equal to the dimension as we prove in Theorem 1. Thus, the proof is completed. ∎
Next, we will prove the upper bound of code distance of 4D XYZ product code in corollary 4. Before proving it, we give out Lemma 2 which will be used in the proof of corollary 4.
Lemma 2.
Giving a vector space $\mathcal{C}_{n}$ with dimension $n$ over $GF(2)$, two binary matrices $H_{1}$ and $H_{2}$ with size $n\times m_{1}$ and $m_{2}\times n$ respectively, the minimum weight of non-zero vector in $\mathcal{C}_{n}\backslash Im(H_{1})$ and $\mathcal{C}_{n}\backslash\ker(H_{2})$ is one.
The proof of Lemma 2 is given in the Appendix A.
Corollary 4 (The upper bound of code distance of 4D XYZ product code).
The upper bound of code distance of 4D XYZ product code is
$d\leq\min\{d_{1},d_{2},d_{3},d_{4}\}$ (42)
where $d_{1}$, $d_{2}$, $d_{3}$ and $d_{4}$ is the minimum weight of the non-zero vectors in spaces $\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)$, $\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, $\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)$ and $\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$, respectively.
Proof.
As proved in Theorem 2, the logical operators of 4D XYZ product code can be divided into two types. Thus, we should consider the upper bound of the minimum weight of these two types of logical operators separately.
First considering the first type of logical operators, and let $\alpha$ and $\beta$ (whose weight is $d_{\beta}$) be minimal weight non-zero vectors in the spaces $\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)$ and $\left(\mathcal{C}_{n_{B}}\backslash Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$ respectively, then we have $\alpha\otimes\beta\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\left(\mathcal{C}_{n_{B}}\backslash Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$ and the weight of $\alpha\otimes\beta$ is $d_{1}d_{\beta}$. According to Lemma 2, we have $d_{\beta}=1$. Thus the weight of $\alpha\otimes\beta$ is $d_{1}$. For $\left(\mathcal{C}_{n_{A}}\backslash Im\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, we have similar conclusion. Thus the upper bound of minimum weight of the vectors in space $\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\left(\mathcal{C}_{n_{B}}\backslash Im\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$ and $\left(\mathcal{C}_{n_{A}}\backslash Im\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is $d_{1}$ and $d_{2}$, respectively.
Secondly, considering the second type of logical operators, and the upper bound of the minimum weight of vector ${\pi\left(\textbf{\emph{a}}_{1}^{T},\textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is $d_{3}$. Observing Eq. (30), it is obvious that the minimum weight of the logical operators $X_{L_{2}}$ and $X_{L_{3}}$ is no more than $d_{3}$. Similarly, the the minimum weight of the logical operators $Z_{L_{2}}$ and $Z_{L_{3}}$ is no more than $d_{4}$.
To sum up, the upper bound of code distance is no more than the minimum of $d_{1}$, $d_{2}$, $d_{3}$ and $d_{4}$. ∎
Here we exploit Monte Carlo method proposed in Ref. [19, 20] to determine code distance of 4D Chamon codes and 4D XYZ concatenated codes and the results are consistent with Corollary 4. Further, we compare the dimension and code distance between 4D Chamon codes, 4D toric codes, 4D XYZ concatenated codes and 4D homological concatenated codes in Table 1, where $n_{1},n_{2},n_{3},n_{4}$ are the code length of the repetition codes. This comparison results support that using the same two CSS codes, the non-CSS code constructed by 4D XYZ product may have higher dimension or code distance than those of the CSS code constructed by 4D homological product.
Table 1: Comparison of 4D Chamon codes, 4D toric codes, 4D XYZ concatenated codes and 4D homological concatenated codes in dimension and code distance.
$n_{1}$, $n_{2}$, $n_{3}$, $n_{4}$ dimension code distance
4D Chamon codes $2,2,2,2$ 32 4
$3,3,3,3$ 72 6
$4,4,4,4$ 128 8
$5,5,5,5$ 200 20
$2,3,2,3$ 8 6
$3,4,3,4$ 8 12
$4,5,4,5$ 8 20
4D toric codes $2,2,2,2$ 6 4
$3,3,3,3$ 6 9
$4,4,4,4$ 6 16
$5,5,5,5$ 6 25
$2,3,2,3$ 6 4
$3,4,3,4$ 6 9
$4,5,4,5$ 6 16
4D XYZ $3,3,3,3$ 1 9
concatenated codes $5,5,5,5$ 1 25
$7,7,7,7$ 1 49
$3,5,3,5$ 1 15
$3,7,3,7$ 1 21
4D homological $2,2,2,2$ 1 4
concatenated codes $5,5,5,5$ 1 25
$7,7,7,7$ 1 49
$3,5,3,5$ 1 9
$3,7,3,7$ 1 9
IV Numerical simulations
In this section, we exploit FDBP-OSD proposed in Ref. [16] to study the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes under the independent single-qubit Pauli noise model, in which each qubit independently suffers a Pauli error $I$, $X$, $Y$, or $Z$ with error probability $1-p$, $p_{x}$, $p_{y}$, $p_{z}$, respectively, where $p=p_{x}+p_{y}+p_{z}$ is physical qubit error rate. Under this noise model, the bias rate of Pauli $Z$ noise is defined as
$\eta=\frac{p_{z}}{p_{x}+p_{y}}$ (43)
For instance, the bias rate of depolarizing and pure Pauli $Z$ noise model is $\eta=0.5$ and $\eta=\infty$, respectively. In this paper, we only consider noiseless stabilizer measurements.
We should emphasize that, in this paper we do not explore an efficient and highly accurate decoder for 3D and 4D XYZ product codes, and FDBP we used here is an improved binary BP, which can be applied to all quantum stabilizer codes and have better decoding performance than tradition binary BP[21, 22], especially for some non-CSS codes[16]. However, as shown in Ref. [23], even though tradition binary BP-OSD[24] has been shown to have pretty good decoding performance for a wide range of QLDPC codes, but its performance can degrade for the codes with large code length and high degeneracy. FDBP also has similar problem and its decoding performance is also greatly degraded by short cycles, especially 4-cycles in the Tanner graph. In Appendix B, we compare the number of 4-cycles in the Tanner graphs of 3D Chamon code, 3D toric codes, 4D Chamon codes and 4D toric codes. Our computed result shows that the number of 4-cycles in the Tanner graph of 3D XYZ and 4D XYZ product codes is much higher than that of 3D and 4D homological product codes. Thus, the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes obtained by FDBP-OSD in this paper is not the optimal. Nonetheless, from the perspective of code-capacity threshold, they still show good error-correcting performance for $Z$-biased noise model.
IV.1 3D Chamon code
3D Chamon code proposed in Ref. [14] can be seen as 3D XYZ product of three repetition codes[9]. Bravyi et al. study some characteristics of this code in depth in Ref. [15]. However, its error-correcting performance has not been studied in depth so far.
The dimension of 3D Chamon code is $4\gcd{\left(n_{1},n_{2},n_{3}\right)}$, where $n_{1}$, $n_{2}$ and $n_{3}$ are the code length of three repetition codes, respectively. Besides, we can also understand 3D Chamon code from the perspective of three-dimensional layout. As shown in Fig. 6, on a three-dimensional cubic Bravais lattice, the qubits of 3D Chamon code are placed on the face centers each of which neighbors two cubes and on the vertices each of which is shared in eight cubes. Each edge of cube represents a stabilizer that involves neighboring six qubits. These six qubits are divided into three classes, each of which contains two qubits along one of the $x$, $y$, $z$ three directions. For each class of qubits, the type of the operators that the stabilizer acts on them is Pauli X or Pauli $Y$ or Pauli $Z$ according to the type of direction. We conjecture that 3D Chamon code may have the same error-correcting capability for Pauli $X$, $Z$ and $Y$ errors when $n_{1}=n_{2}=n_{3}$, since there exists spatial symmetry in this case.
Figure 6: Three-dimensional layout of Chamon code. Qubits are placed on the face centers and the vertices of the cubes. Each edge of the cube represents a stabilizer, which involves neighboring six qubits. Some qubits are omitted for clearly.
[…]
Figure 7: The error-correcting performance of Chamon codes under (a) pure Pauli $X$ noise model, (b) pure Pauli $Z$ noise model, (c) pure Pauli $Y$ noise model and (d) depolarizing noise model.
IV.2 4D Chamon code
We name the code which is 4D XYZ product of four repetition codes with code length $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ the 4D Chamon code and conjecture that it is 4D generalization of 3D Chamon code, just as the relation between 3D toric code and 4D toric code. There are two reasons. First, from the perspective of code dimension, the dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, which is similar with $4\gcd{\left(n_{1},n_{2},n_{3}\right)}$, the dimension of 3D Chamon code. Second, from the perspective of code construction, the tensor structures of 3D and 4D XYZ product, as shown in Fig. 2 and 3, are the same as those of 3D and 4D homological product, respectively, while 3D and 4D toric codes are constructed by 3D and 4D homological product, respectively. We also conjecture that 3D and 4D Chamon codes may have similar topological structures, but in this paper we do not make further exploration for it.
In Fig. 8, we first consider the case of $n_{1}=n_{2}=n_{3}=n_{4}$. Fig. 8(a) and (b) show the single logical qubit error rate against the physical qubit error rate $p$ of 4D Chamon codes and 4D toric codes under depolarizing noise model, respectively. In this case, the 4D Chamon code encodes $8n_{1}^{2}$ logical qubits, while the 4D toric codes always encodes 6 logical qubits. In the legend, we use notation $L=n_{1},n_{2},n_{3},n_{4}$ to represent the corresponding the code, which is constructed by four repetition codes with code length of $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$. One can see that the code-capacity threshold of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD is around $12.8\%$ and $15.9\%$, respectively, and the single logical qubit error rate of 4D Chamon codes is much higher than that of 4D toric codes. We attribute this to the fact that the number of 4-cycles of 4D Chamon codes is much more than that of 4D toric codes as shown in Appendix B, which degrades the decoding performance of FDBP-OSD.
Figure 8: When $n_{1}=n_{2}=n_{3}=n_{4}$, the error-correcting performance of (a) 4D Chamon codes and (b) 4D toric codes under depolarizing noise model.
The dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, when $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the corresponding code always encode 8 logical qubits, which is close to 4D toric codes. Thus, it is more reasonable to compare the error-correcting performance between 4D Chamon codes and 4D toric codes in this case. First, we consider the pure Pauli $Z$ noise model. As shown in Fig. 9 (a) and (b), the code capacity of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD under Pauli $Z$ noise is around $18\%$ and $9.5\%$, respectively. This result means, form the perspective of code-capacity threshold, 4D Chamon codes have better error-correcting performance than that of 4D toric codes under pure Pauli $Z$ noise.
Figure 9: When $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the error-correcting performance of (a) 4D Chamon codes and (b) 4D toric codes under pure Pauli $Z$ noise model.
Further, we explore the error-correcting performance of 4D Chamon codes under $Z$-biased noise with bias rate $\eta=1000$, 100, 10 and 0.5. Our simulation results show that the corresponding code-capacity threshold is around $19\%$, $18\%$, $17\%$ and $13\%$, respectively.
Figure 10: When $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, The error-correcting performance of 4D Chamon codes under $Z$-biased noise with bias rate (a) $\eta=1000$, (b) 100, (c) 10 and (d) 0.5.
IV.3 4D XYZ concatenated code
As described in Sect. III.2, 4D XYZ concatenated code is constructed from two concatenated code, which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ ($n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and only encodes one logical qubit. As proved in Sect. III.3 the minimum weight of its logical $Z$ operator is $n_{3}n_{4}$, thus we conjecture that increasing block lengths $\left(n_{3},\ n_{4}\right)$ of the pair of repetition codes can construct the corresponding 4D XYZ product codes, which have better error-correcting performance for $Z$-biased noise than the codes constructed by 4D homological product.
[…]
Figure 11: The error-correcting performance of (a) 4D XYZ concatenated code and (b) 4D homological product codes under pure $Z$ noise model.
Further, we explore the error-correcting performance of 4D XYZ concatenated codes under Z-biased noise with bias rate $\eta=1000$, 100, 10 and 0.5. Our simulation results show that the corresponding code-capacity threshold is around $37\%$, $32\%$, $22\%$ and $10\%$, respectively, which support that 4D XYZ product code can construct codes with good error-correcting performance for $Z$-biased noise.
Figure 12: The error-correcting performance of 4D XYZ concatenated codes under Z-biased noise with bias rate (a) $\eta=1000$, (b) 100, (c) 10 and (d) 0.5.
V Conclusion
[…]
The results above mean that using the same two CSS codes, 4D XYZ product can construct non-CSS codes with better error-correcting performance for $Z$-biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
Appendix A Proof of Lemma 2
Proof.
Here, we prove the minimum weight of non-zero vectors in space $\mathcal{C}_{n}\backslash Im(H_{1})$ with that of $\mathcal{C}_{n}\backslash\ker(H_{2})$ following a similar fashion.
The space $\mathcal{C}_{n}$ can be represented by the row space of an $n\times n$ identity matrix $I_{n}$, namely, $row(I_{n})$. Thus, this problem is equivalent to finding the minimum weight of non-zero vectors in space $row(I_{n})$ but not in $Im(H_{1})$. To do this, one can construct a block matrix $\left[H_{1},I_{n}\right]$, then using Gaussian elimination to find out a set of column indices of its pivots P. For the rows which are in $I_{n}$ but not in $H_{1}^{T}$ and whose indices are in P, they are the vectors in space $row(I_{n})$ but not in $Im(H_{1})$. Since the weight of each row of $I_{n}$ is one, thus the minimum weight of non-zero vectors in space $\mathcal{C}_{n}\backslash Im(H_{1})$ is one and the proof is completed. ∎
Appendix B Comparison of number of 4-cycles
Table 2 gives the number of 4-cycles in the Tanner graphs of 3D Chamon codes, 4D Chamon codes, 3D toric codes and 4D toric codes.
Table 2: Comparison of 3D Chamon codes, 4D Chamon codes, 3D toric codes and 4D toric codes in the number of 4-cycles.
$n_{1}$, $n_{2}$, $n_{3}$ number of 4-cycles
3D Chamon codes $2,2,2$ 240
$3,3,3$ 648
$4,4,4$ 1536
$2,3,4$ 624
$3,4,5$ 1440
3D toric codes $2,2,2$ 132
$3,3,3$ 324
$4,4,4$ 768
$2,3,4$ 201
$3,4,5$ 467
$n_{1}$, $n_{2}$, $n_{3}$, $n_{4}$ number of 4-cycles
4D Chamon codes $2,2,2,2$ 1792
$2,3,2,3$ 3744
4D toric codes $2,2,2,2$ 519
$2,3,2,3$ 1095
End Notes
Acknowledgements
This work was supported by the Colleges and Universities Stable Support Project of Shenzhen, China (No.GXWD20220817164856008), the Colleges and Universities Stable Support Project of Shenzhen, China (No.GXWD20220811170225001) and Harbin Institute of Technology, Shenzhen - SpinQ quantum information Joint Research Center Project (No.HITSZ20230111).
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.